OpenStudy (anonymous):

Simplify the expression. square root of 128 over square root of 2

OpenStudy (dwelsher):

64

OpenStudy (anonymous):

\[\sqrt(128)=\sqrt(2^7)\] \[\sqrt(2^7)/\sqrt(2)=\sqrt(2^6)=2^3=8\]

Directrix (directrix):

@dwelsher Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.